OpenStudy (saad560):

What is the value of f? 5f+18=4f−2 −20 −16 16 20

OpenStudy (3mar):

First: you need to bring the variable terms in one side and others in the other side.

OpenStudy (mathmale):

Group the f terms on the left side of your equation and the constants on the right. What do you get? Now, simplify both sides.

OpenStudy (3mar):

Recalculate, please. It is easy.

OpenStudy (mathmale):

Hint: SUBTRACT 4f from both sides. Share your result, please.

OpenStudy (saad560):

@563blackghost

OpenStudy (3mar):

Can you "SUBTRACT 4f from both sides"?

563blackghost (563blackghost):

@3mar and @mathmale are providing great help :)

563blackghost (563blackghost):

What they are saying is to get the variable to one side....so they are telling you to subtract 4f from both sides... \(\huge{(5f-4f)+18=(4f-4f)-2}\)

OpenStudy (saad560):

1f+18=-2

OpenStudy (mathmale):

Great. You have isolated "f" on the left side. Now you must regroup so that all the constant terms are on the right side. How would you do that? Leave the 1f alone.

OpenStudy (saad560):

i am confused

OpenStudy (saad560):

@563blackghost

563blackghost (563blackghost):

What @mathmale is saying is that yes you are correct so the equation is now... \(\huge{f+18=-2}\) Now he wants you to say which is the next step....we need to isolate f (put it to itself) so what would we do??

OpenStudy (mathmale):

Do you want 18 to stay on the left side of this equation? Yes or no? If no, decide how to get rid of the 18. See what 563Blackghost told you earlier: "What they are saying is to get the variable to one side....so they are telling you to subtract 4f from both sides..."

563blackghost (563blackghost):

We have... \(\huge{f+18=-2}\) Now we need to get f as... \(\huge{f=}\) This is isolating it....since we have f is added to 18 in order to cancel it out we do the opposite of the sign of 18....

OpenStudy (3mar):

\[f+18=-2\] \[f+18-18=-2-18\] \[f+0=-20\] \[f=-20\] Done!